English

The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1.

Advertisements
Advertisements

Question

The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English Examination is 0.75. What is the probability of passing the Hindi Examination?

Sum
Advertisements

Solution

Let A and B be the events of passing English and Hindi examinations, respectively.

Accordingly, we have:

P(A and B) = 0.5

 P(not A and not B) = 0.1 [i.e. P(A' ∩ B') = 0.1]

P(A) = 0.75

Now, P(A∪B) + P(A' ∩ B') = 1

⇒ P(A∪B) = 1 - P(A' ∩ B')

                  = 1 -0.1 = 0.9

By addition theorem, we have:

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

⇒ 0.9 = 0.75 + P (B)  - 0.5

⇒ P(B) = 0.9 - 0.75 + 0.5

⇒ P(B) = 0.65

Thus, the probability of passing the Hindi examination is 0.65.

shaalaa.com
Concept of Probability - Probability of 'Not', 'And' and 'Or' Events
  Is there an error in this question or solution?
Chapter 14: Probability - EXERCISE 14.2 [Page 307]

APPEARS IN

NCERT Mathematics [English] Class 11
Chapter 14 Probability
EXERCISE 14.2 | Q 20. | Page 307
R.D. Sharma Mathematics [English] Class 11
Chapter 33 Probability
Exercise 33.4 | Q 13 | Page 68

RELATED QUESTIONS

A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is

  1. a vowel
  2. an consonant

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(not A).


A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P (not B)


In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?


The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?


In a certain lottery, 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy one ticket.


Out of 100 students, two sections of 40 and 60 are formed. If you and your friend are among the 100 students, what is the probability that

  1. you both enter the same sections?
  2. you both enter the different sections?

A dice is thrown. Find the probability of getting a prime number


In a simultaneous throw of a pair of dice, find the probability of getting a doublet


In a simultaneous throw of a pair of dice, find the probability of getting  an even number on first


In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 6


In a simultaneous throw of a pair of dice, find the probability of getting a sum less than 7


In a simultaneous throw of a pair of dice, find the probability of getting a number greater than 4 on each die


In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.

 

Two dice are thrown. Find the odds in favour of getting the sum 5.

 

 


Two dice are thrown. Find the odds in favour of getting the sum  What are the odds against getting the sum 6?


What are the odds in favour of getting a spade if the card drawn from a well-shuffled deck of cards? What are the odds in favour of getting a king?

 

A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn at random. From the box, what is the probability that  at least one is green?


A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.


If A and B are two events associated with a random experiment such that
P (A ∪ B) = 0.8, P (A ∩ B) = 0.3 and P \[(\bar{A} )\]= 0.5, find P(B).

 


A card is drawn at random from a well-shuffled deck of 52 cards. Find the probability of its being a spade or a king.


A card is drawn from a deck of 52 cards. Find the probability of getting an ace or a spade card.


100 students appeared for two examination, 60 passed the first, 50 passed the second and 30 passed both. Find the probability that a student selected at random has passed at least one examination.


A person write 4 letters and addresses 4 envelopes. If the letters are placed in the envelopes at random, then the probability that all letters are not placed in the right envelopes, is


If the probability of A to fail in an examination is \[\frac{1}{5}\]  and that of B is \[\frac{3}{10}\] . Then, the probability that either A or B fails is

 
 

A box contains  10 good articles and 6 defective articles. One item is drawn at random. The probability that it is either good or has a defect, is


Out of 30 consecutive integers, 2 are chosen at random. The probability that their sum is odd, is


A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is black or red ball is


Two dice are thrown simultaneously. The probability of getting a pair of aces is


An urn contains 9 balls two of which are red, three blue and four black. Three balls are drawn at random. The probability that they are of the same colour is


Five persons entered the lift cabin on the ground floor of an 8 floor house. Suppose that each of them independently and with equal probability can leave the cabin at any floor beginning with the first, then the probability of all 5 persons leaving at different floor is


A box contains 10 good articles and 6 with defects. One item is drawn at random. The probability that it is either good or has a defect is


One mapping is selected at random from all mappings of the set A = {1, 2, 3, ..., n} into itself. The probability that the mapping selected is one to one is


In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy 10 tickets.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×